# Exploratory Randomization for Discrete-Time Linear Exponential Quadratic Gaussian (LEQG) Problem

### v1

@@ Abstract
We investigate exploratory randomization for an extended linear-exponential-quadratic-Gaussian (LEQG) control problem in discrete time. This extended control problem is related to the structure of risk-sensitive investment management applications. We introduce exploration through a randomization of the control. Next, we apply the duality between free energy and relative entropy to reduce the LEQG problem to an equivalent risk-neutral LQG control problem with an entropy regularization term, see, e.g. Dai Pra et al., (1996), for which we present a solution approach based on Dynamic Programming. Our approach, based on the energy-entropy duality may also be considered as leading to a justification for the use, in the literature, of an entropy regularization when applying a randomized control.
@@ 1 Introduction
Real-world stochastic control problems are often characterized by incomplete information about the state or the model. Traditionally, filtering has been used to address incomplete information (see Bensoussan & Van Schuppen,, 1984; Bensoussan & Runggaldier,, 1987). More recently, reinforcement learning has emerged as a promising alternative (see Wang et al.,, 2020; Jia & Zhou,, 2023; Jia & Zhou, 2022a, ; Jia & Zhou, 2022b, ; Hambly et al.,, 2021). These techniques rely on iterative interactions with the environment, balancing exploration—randomized actions to acquire information—and exploitation—actions that maximize a reward function based on available data. Although reinforcement learning evolved from Markov Decision Processes, integrating it with stochastic control remains challenging. Wang et al. (2020) made a key advance in this direction by showing how relaxed controls create exploratory policies inside standard (risk-neutral) control problems formulated in continuous time. Adding an entropy penalization/regularization to the control criterion further led to a trade-off between exploration and exploitation.
@@
A particularly important class of stochastic control problems is that of risk-sensitive control (see Whittle,, 1990; Bensoussan et al.,, 1998; Bielecki et al.,, 2022; Bäuerle & Rieder,, 2014; Bäuerle & Jaśkiewicz,, 2024), with linear exponential of quadratic Gaussian (LEQG) problems as a notable subclass (see Jacobson,, 1973; Whittle,, 1981; Bensoussan & van Schuppen,, 1985; Bensoussan,, 1992). LEQG problems are especially relevant in financial applications (see e.g. Bielecki & Pliska,, 1999; Kuroda & Nagai,, 2002; Davis & Lleo,, 2008, 2014; Pitera & Stettner,, 2023; Davis & Lleo,, 2020, 2021; Lleo & Runggaldier,, 2024, and references therein).
@@
In this paper, we consider a discrete-time risk-sensitive control problem of the LEQG type drawing inspiration from the results of Wang et al., (2020). Specifically, we introduce randomized controls to emulate exploration. We generate these randomized controls by perturbing a deterministic control with an additive Gaussian distribution. Here, the mean is the deterministic control that captures exploitation, while the variance, which provides an additional (antagonizing) control, models exploration. While implementing additive randomization in continuous-time settings poses significant technical challenges [^1] It is worth noting that alternative continuous-time approaches, such as those developed by Huyên Pham and coauthors (for a recent contribution see e.g. Denkert et al., (2024)), provide valuable insights., the discrete-time framework allows us to focus more directly on the exploration aspect.
@@
Our main contribution is to solve the randomized LEQG problem using a duality between free energy and entropy Dai Pra et al., (1996). This approach reduces the risk-sensitive randomized stochastic control problem to a risk-neutral stochastic game problem with an entropy penalization. This penalization comes from two sources: (i) the randomized controls and (ii) the transition from risk-sensitive to risk-neutral. While Wang et al., (2020) introduce an entropy penalization term as a regularisation, the entropy penalization in our approach arises naturally from the energy-entropy duality, providing a conceptual justification for its use. Unlike Wang et al., who remain in the domain of standard risk-neutral stochastic control problems, our formulation yields a stochastic game. In this game, the decision-maker selects controls to exploit available information, while an antagonistic player determines the variance of the control perturbations, governing the degree of exploration.
@@
The introduction of randomization into stochastic control problems is primarily motivated by its applicability in reinforcement learning. A possible methodology to implement reinforcement learning in our setup is via a policy gradient approach, whereby learning is applied to a parametrized optimal policy (see for instance Hambly et al.,, 2021). However, our paper focuses on the foundational modeling steps that precede the development of the reinforcement learning algorithm. Specifically, we demonstrate how to reduce the randomized LEQG problem to a randomized stochastic LQG game via the energy-entropy duality. We also derive the optimal policy and value function. Both turn out to be parametrized by the initial data of the model via a backward recursion, as stated in Theorem 3.5. If the model parameters are updated by some method that we do not explore here, the backward recursion has to be recomputed. In this sense, our approach retains conceptual similarities with the policy gradient framework, and in particular actor-critic methods.
@@ 2.1 Linear Exponential Quadratic (LEQG) Control Problem
The risk-sensitive criterion $J$ to be minimized is defined as
@@
where $\theta\in(-1,0)\cup(0,\infty)$ is the risk-sensitivity,
@@
$M\in\mathbb{R}^{d_{x}\times d_{x}}$ is semipositive definite and symmetric, $N_{t}\in\mathbb{R}^{d_{u}\times d_{u}},t=1,\ldots,T-1$ is also semipositive definite and symmetric, $Q\in\mathbb{R}^{d_{u}\times d_{x}}$, $m\in\mathbb{R}^{d_{x}}$, $n\in\mathbb{R}^{d_{u}}$, $M_{T}\in\mathbb{R}^{d_{x}\times d_{x}}$, and $m_{T}\in\mathbb{R}^{d_{x}}$. For convenience, we also introduce the exponentially transformed criterion $I$ defined as
$I(u;T,\theta)=e^{-\theta J(u;T,\theta)}=\mathbf{E}\left[e^{\theta G_{T}}\right].$ (2.4)
@@
The standard formulation of risk-sensitive control problems does not contain a constant $a$ in the state dynamics or a cross-term $u_{t}Qx_{t}$ in the running reward. These additions are required in investment management applications of continuous-time risk-sensitive control, see for example Davis & Lleo, (2014). Including them ensures that our model can be used to solve discretized versions of the risk-sensitive asset management model by Bielecki & Pliska, (1999); Kuroda & Nagai, (2002) and of the risk-sensitive benchmarked asset management model by Davis & Lleo, (2008).
@@ 2.2 Exploration via Randomized Controls
The classical LEQG model may not provide a perfect description of reality, and we may not have sufficient data to estimate our model parameters with a high degree of accuracy. We employ exploratory randomized controls to address these issues. Wang et al., (2020) already proposed using randomized controls in continuous-time stochastic control problems to create the kind of exploration framework that is already customary in reinforcement learning methods. These authors showed that under suitable assumptions, adding an entropy-related penalization to their stochastic control framework produces an optimal exploration-exploitation trade-off.
@@
Taking into consideration the randomization of controls, we express the elements $\omega$ of the underlying sample space $\Omega$ as $\omega=\left(\omega^{w},\omega^{v}\right):=\left(w_{0},\ldots,w_{T-1},v_{0},\ldots,v_{T-1}\right)$.
@@
for $t=0,\ldots,T-1$. Inspired by Wang et al., we express the reward as:
$G_{T}:=$ $\sum_{t=0}^{T-1}\left[x_{t}^{\prime}Mx_{t}+\int u_{t}^{\prime}N_{t}u_{t}\pi\left(du;\bar{u}_{t}\right)+\int u_{t}^{\prime}Qx_{t}\pi\left(du;\bar{u}_{t}\right)+x_{t}^{\prime}m+\int u_{t}^{\prime}n\pi\left(du;\bar{u}_{t}\right)\right]$
$+x_{T}^{\prime}M_{T}x_{T}+x_{T}^{\prime}m_{T}$
::=:= $\sum_{t=0}^{T-1}\left[x_{t}^{\prime}Mx_{t}+\text{tr}\left(\Xi_{t}N_{t}\right)+\bar{u}_{t}^{\prime}N_{t}\bar{u}_{t}+\bar{u}_{t}^{\prime}Qx_{t}+x_{t}^{\prime}m+\bar{u}_{t}^{\prime}n\right]+x_{T}^{\prime}M_{T}x_{T}+x_{T}^{\prime}m_{T}.$ (2.6)
@@ 3 Solving the Randomized LEQG Control Problem via the Free Energy-Entropy Duality
We start this section by recalling key results from Dai Pra et al., (1996) about the duality between free energy and entropy. We shall next use the energy-entropy duality to associate to the given randomized LEQG problem a risk-neutral randomized stochastic game with a penalization given by an appropriate relative entropy. This penalization acts on both the transition from a risk-sensitive to risk-neutral problem and the randomization of controls.
@@ 3.1 General Definitions
[…]
@@
where $\mathbf{E}^{\gamma}\left[\cdot\right]$ denotes the expectation with respect to the measure $\mathbb{P}^{\gamma}$. The free energy-entropy duality relation (see (ii) in Proposition 2.3. in Dai Pra et al.,, 1996) is
[…]
@@
Dai Pra et al. also provide an analytical formula that the optimal Radon-Nikodym derivative $\frac{d\mathbb{P}^{\gamma^{*}}}{d\mathbb{P}}\Big{|}_{\mathcal{F}_{t}}$ must satisfy:
$\frac{d\mathbb{P}^{\gamma^{*}}}{d\mathbb{P}}\Big{|}_{\mathcal{F}_{t}}=\frac{e^{\Psi}}{\mathbf{E}^{\mathbb{P}}\left[e^{\Psi}\right]},$ (3.4)
under suitable conditions presented in Dai Pra et al., (1996).

### v2

@@ Abstract
We investigate exploratory randomization for an extended linear-exponential-quadratic-Gaussian (LEQG) control problem in discrete time. This extended control problem is related to the structure of risk-sensitive investment management applications. We introduce exploration through a randomization of the control. Next, we apply the duality between free energy and relative entropy to reduce the LEQG problem to an equivalent risk-neutral LQG control problem with an entropy regularization term, see, e.g. Dai Pra et al. (1996), for which we present a solution approach based on Dynamic Programming. Our approach, based on the energy-entropy duality may also be considered as leading to a justification for the use, in the literature, of an entropy regularization when applying a randomized control.
keywords: risk-sensitive control, free energy-entropy duality, stochastic differential games, Linear Exponential Quadratic Gaussian control, Linear Quadratic Gaussian games, exploratory controls.
@@ 1 Introduction
Real-world stochastic control problems are often characterized by incomplete information about the state or the model. Traditionally, filtering has been used to address incomplete information (see Bensoussan and Van Schuppen, 1984; Bensoussan and Runggaldier, 1987). More recently, reinforcement learning has emerged as a promising alternative (see Wang et al., 2020; Jia and Zhou, 2023, 2022a, 2022b; Hambly et al., 2021). These techniques rely on iterative interactions with the environment, balancing exploration—randomized actions to acquire information—and exploitation—actions that maximize a reward function based on available data. Although reinforcement learning evolved from Markov Decision Processes, integrating it with stochastic control remains challenging. Wang et al. (2020) made a key advance in this direction by showing how relaxed controls create exploratory policies inside standard (risk-neutral) control problems formulated in continuous time. Adding an entropy penalization/regularization to the control criterion further led to a trade-off between exploration and exploitation.
@@
A particularly important class of stochastic control problems is that of risk-sensitive control (see Whittle, 1990; Bensoussan et al., 1998; Bielecki et al., 2022; Bäuerle and Rieder, 2014; Bäuerle and Jaśkiewicz, 2024), with linear exponential of quadratic Gaussian (LEQG) problems as a notable subclass (see Jacobson, 1973; Whittle, 1981; Bensoussan and van Schuppen, 1985; Bensoussan, 1992). LEQG problems are especially relevant in financial applications (see e.g. Bielecki and Pliska, 1999; Kuroda and Nagai, 2002; Davis and Lleo, 2008, 2014; Pitera and Stettner, 2023; Davis and Lleo, 2020, 2021; Lleo and Runggaldier, 2024, and references therein).
@@
In this paper, we consider a discrete-time risk-sensitive control problem of the LEQG type drawing inspiration from the results of Wang et al. (2020). Specifically, we introduce randomized controls to emulate exploration. We generate these randomized controls by perturbing a deterministic control with an additive Gaussian distribution. Here, the mean is the deterministic control that captures exploitation, while the variance, which provides an additional (antagonizing) control, models exploration. While implementing additive randomization in continuous-time settings poses significant technical challenges [^1] It is worth noting that alternative continuous-time approaches, such as those developed by Huyên Pham and coauthors (for a recent contribution see e.g. Denkert et al. (2024)), provide valuable insights., the discrete-time framework allows us to focus more directly on the exploration aspect.
@@
Our main contribution is to solve the randomized LEQG problem using a duality between free energy and entropy Dai Pra et al. (1996). This approach reduces the risk-sensitive randomized stochastic control problem to a risk-neutral stochastic game problem with an entropy penalization. This penalization comes from two sources: (i) the randomized controls and (ii) the transition from risk-sensitive to risk-neutral. While Wang et al. (2020) introduce an entropy penalization term as a regularisation, the entropy penalization in our approach arises naturally from the energy-entropy duality, providing a conceptual justification for its use. Unlike Wang et al., who remain in the domain of standard risk-neutral stochastic control problems, our formulation yields a stochastic game. In this game, the decision-maker selects controls to exploit available information, while an antagonistic player determines the variance of the control perturbations, governing the degree of exploration.
@@
The introduction of randomization into stochastic control problems is primarily motivated by its applicability in reinforcement learning. A possible methodology to implement reinforcement learning in our setup is via a policy gradient approach, whereby learning is applied to a parametrized optimal policy (see for instance Hambly et al., 2021). However, our paper focuses on the foundational modeling steps that precede the development of the reinforcement learning algorithm. Specifically, we demonstrate how to reduce the randomized LEQG problem to a randomized stochastic LQG game via the energy-entropy duality. We also derive the optimal policy and value function. Both turn out to be parametrized by the initial data of the model via a backward recursion, as stated in Theorem 3.5. If the model parameters are unknown and are updated recursively by some method that we only sketch briefly, the backward recursion has to be recomputed. In this sense, our approach retains conceptual similarities with the policy gradient framework, and in particular, actor-critic methods as further illustrated in Section 4.3 below.
@@ 2.1 Linear Exponential Quadratic (LEQG) Control Problem
The risk-sensitive criterion $J$ to be maximized is defined as
@@
where $\theta\in(-1,0)\cup(0,\infty)$ is the risk-sensitivity, $G_{T}$ is the cost defined as
@@
where $M\in\mathbb{R}^{d_{x}\times d_{x}}$ is semipositive definite and symmetric, $N_{t}\in\mathbb{R}^{d_{u}\times d_{u}},t=1,\ldots,T-1$ is also semipositive definite and symmetric, $Q\in\mathbb{R}^{d_{u}\times d_{x}}$, $m\in\mathbb{R}^{d_{x}}$, $n\in\mathbb{R}^{d_{u}}$, $M_{T}\in\mathbb{R}^{d_{x}\times d_{x}}$, and $m_{T}\in\mathbb{R}^{d_{x}}$. For convenience, we also introduce the exponentially transformed criterion $I$ defined as
$I(u;T,\theta)=e^{-\theta J(u;T,\theta)}=\mathbf{E}\left[e^{\theta G_{T}}\right],$ (2.4)
which is to be minimized.
@@
The standard formulation of risk-sensitive control problems does not contain a constant $a$ in the state dynamics or a cross-term $u_{t}Qx_{t}$ in the running cost. These additions are required in investment management applications of continuous-time risk-sensitive control, see for example Davis and Lleo (2014). Including them ensures that our model can be used to solve discretized versions of the risk-sensitive asset management model by Bielecki and Pliska (1999); Kuroda and Nagai (2002) and of the risk-sensitive benchmarked asset management model by Davis and Lleo (2008).
@@ 2.2 Exploration via Randomized Controls
The classical LEQG model may not provide a perfect description of reality, and we may not have sufficient data to estimate our model parameters with a high degree of accuracy. We employ exploratory randomized controls to address these issues. Wang et al. (2020) already proposed using randomized controls in continuous-time stochastic control problems to create the kind of exploration framework that is already customary in reinforcement learning methods. These authors showed that under suitable assumptions, adding an entropy-related penalization to their stochastic control framework produces an optimal exploration-exploitation trade-off.
@@
Taking into consideration the randomization of controls, we express the elements $\omega$ of the underlying sample space $\Omega$ as
$\omega=\left(\omega^{w},\omega^{v}\right):=\left(w_{0},\ldots,w_{T-1},v_{0},\ldots,v_{T-1}\right).$
@@
for $t=0,\ldots,T-1$. Inspired by Wang et al., we express the cost as:
$G_{T}:=$ $\sum_{t=0}^{T-1}\left[x_{t}^{\prime}Mx_{t}+\int u_{t}^{\prime}N_{t}u_{t}\pi\left(du;\bar{u}_{t}\right)+\int u_{t}^{\prime}Qx_{t}\pi\left(du;\bar{u}_{t}\right)\right$. (2.6)
$\left.+x_{t}^{\prime}m+\int u_{t}^{\prime}n\pi\left(du;\bar{u}_{t}\right)\right]+x_{T}^{\prime}M_{T}x_{T}+x_{T}^{\prime}m_{T}$
$=$ $\sum_{t=0}^{T-1}\left[x_{t}^{\prime}Mx_{t}+\text{tr}\left(\Xi_{t}N_{t}\right)+\bar{u}_{t}^{\prime}N_{t}\bar{u}_{t}+\bar{u}_{t}^{\prime}Qx_{t}+x_{t}^{\prime}m+\bar{u}_{t}^{\prime}n\right]$
$+x_{T}^{\prime}M_{T}x_{T}+x_{T}^{\prime}m_{T}.$
@@ 3 Solving the Randomized LEQG Control Problem via the Free Energy-Entropy Duality
We start this section by recalling key results from Dai Pra et al. (1996) about the duality between free energy and entropy. We shall next use the energy-entropy duality to associate to the given randomized LEQG problem a risk-neutral randomized stochastic game with a penalization given by an appropriate relative entropy. This penalization acts on both the transition from a risk-sensitive to risk-neutral problem and the randomization of controls.
As a reference for the reader, Table 1 lists the key variables and parameters used in Sections 2 and 3.
Variable/ Dimension Definition Introduced
Parameter in Section
State and total cost:
$x_{t}$ $\mathbb{R}^{d_{x}}$ State variable 2.1
$G_{T}$ $\mathbb{R}$ Total cost 2.1
Under the measure $\mathbb{P}$: $x_{t+1}=a+Ax_{t}+Bu_{t}+w_{t}$
$G_{T}=\sum_{t=0}^{T-1}\left(x_{t}^{\prime}Mx_{t}+u_{t}^{\prime}N_{t}u_{t}+u_{t}^{\prime}Qx_{t}+x_{t}^{\prime}m+u_{t}^{\prime}n\right)$
[…]
$w_{t}$ $\mathbb{R}^{d_{x}}$ State noise: $w_{t}\sim\mathcal{N}\left(0,\Lambda_{t}\right)$ 2.1
$u_{t}$ $\mathbb{R}^{d_{u}}$ 1. Generic control for the initial LEQG problem 2.1
2. Exploratory control: $u_{t}=\bar{u}_{t}+v_{t}\sim\mathcal{N}\left(\bar{u}_{t},\Xi_{t}\right)$ 2.2
$\bar{u}_{t}$ $\mathbb{R}^{d_{u}}$ Deterministic control 2.2
$v_{t}$ $\mathbb{R}^{d_{u}}$ Exploration noise: 2.2
$v_{t}\sim\mathcal{N}\left(0,\Xi_{t}\right)$
Under the measure $\mathbb{P}^{\bar{\gamma},\bar{\eta}}$: $x_{t+1}=a+Ax_{t}+B\left(\bar{u}_{t}+v^{\bar{\eta}}_{t}\right)+w^{\bar{\gamma}}_{t}$
$G_{T}=\sum_{t=0}^{T-1}\Big[x_{t}^{\prime}Mx_{t}+\text{tr}\left(\Xi_{t}N_{t}\right)+\left(\bar{u}_{t}+\bar{\eta}_{t}\right)^{\prime}N_{t}\left(\bar{u}_{t}+\bar{\eta}_{t}\right)\Big]$
[…]
$\bar{\gamma}$ $\mathbb{R}^{d_{x}}$ Change of measure process (control) 3.2
$\bar{\eta}$ $\mathbb{R}^{d_{u}}$ Change of measure process (control) 3.2
$w^{\bar{\gamma}}_{t}$ $\mathbb{R}^{d_{x}}$ State noise: $w^{\bar{\gamma}}_{t}\sim\mathcal{N}\left(\bar{\gamma}_{t},\Lambda_{t}\right)$ 3.2
$v^{\bar{\eta}}_{t}$ $\mathbb{R}^{d_{x}}$ Exploration noise: $v^{\bar{\eta}}_{t}\sim\mathcal{N}\left(\bar{\eta}_{t},\Xi_{t}\right)$ 3.2
Value function for the stochastic differential game:
$V(T;\theta)$ $\mathbb{R}$ Value of the stochastic differential game: 3.2
$V(T;\theta):=\inf_{\bar{u}}\sup_{\bar{\gamma},\bar{\eta}}\mathbf{E}^{\bar{\gamma},\bar{\eta}}\left[\theta G_{T}-\frac{1}{2}\sum_{t=0}^{T-1}\left(\bar{\gamma}_{t}^{\prime}\Lambda^{-1}_{t}\bar{\gamma}_{t}+\bar{\eta}_{t}^{\prime}\Xi_{t}^{-1}\bar{\eta}_{t}\right)\right]$
$V_{t}(x_{t})$ $\mathbb{R}$ Recursive quadratic expression: $V_{t}(x_{t})=\frac{1}{2}x_{t}^{\prime}P_{t}x_{t}+x_{t}^{\prime}p_{t}+r_{t}$ 3.3
Parameters used in the resolution of the control problem:
$\mathfrak{A}^{(1)}_{t+1}$ $\mathbb{R}^{d_{u}}\times\mathbb{R}^{d_{x}}$ $\mathfrak{A}^{(1)}_{t+1}=B^{\prime}P_{t+1}A+\theta Q$ 3.3
$\mathfrak{a}^{(1)}_{t+1}$ $\mathbb{R}^{d_{u}}$ $\mathfrak{a}^{(1)}_{t+1}=B^{\prime}P_{t+1}a+B^{\prime}p_{t+1}+\theta n$ 3.3
$\mathfrak{A}^{(2)}_{t+1}$ $\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{x}}$ $\mathfrak{A}^{(2)}_{t+1}=P_{t+1}^{\prime}A$ 3.3
$\mathfrak{a}^{(2)}_{t+1}$ $\mathbb{R}^{d_{x}}$ $\mathfrak{a}^{(2)}_{t+1}=P_{t+1}^{\prime}a+p_{t+1}$ 3.3
$\mathfrak{B}^{(1)}_{t+1}$ $\mathbb{R}^{d_{u}}\times\mathbb{R}^{d_{u}}$ $\mathfrak{B}^{(1)}_{t+1}=B^{\prime}P_{t+1}B+2\theta N_{t}$ 3.3
$\mathfrak{B}^{(2)}_{t+1}$ $\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{x}}$ $\mathfrak{B}^{(2)}_{t+1}=-\Lambda^{-1}_{t}+P_{t+1}$ 3.3
$\mathfrak{B}^{(3)}_{t+1}$ $\mathbb{R}^{d_{u}}\times\mathbb{R}^{d_{u}}$ $\mathfrak{B}^{(3)}_{t+1}=-\Xi_{t}^{-1}+\mathfrak{B}^{(1)}_{t+1}=-\Xi_{t}^{-1}+B^{\prime}P_{t+1}B+2\theta N_{t}$ 3.3
$\mathfrak{C}_{t+1}$ $\mathbb{R}^{d_{u}}\times\mathbb{R}^{d_{x}}$ $\mathfrak{C}_{t+1}=B^{\prime}P_{t+1}$ 3.3
$\mathfrak{G}_{t+1}$ $\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{x}}$ $\mathfrak{G}_{t+1}:=\mathfrak{C}_{t+1}^{\prime}\left(\mathfrak{B}^{(1)}_{t+1}\right)^{-1}\mathfrak{C}_{t+1}-\mathfrak{B}^{(2)}_{t+1}$ 3.3
Table 1: Key variables and parameters appearing in Sections 2 and 3
@@ 3.1 General Definitions
$\mathcal{E}^{\mathbb{P}}\left\{\psi\right\}=\ln\left(\int e^{\psi}d\mathbb{P}\right).$ (3.1)
@@
where $\mathbf{E}^{\gamma}\left[\cdot\right]$ denotes the expectation with respect to the measure $\mathbb{P}^{\gamma}$. The free energy-entropy duality relation (see (ii) in Proposition 2.3. in Dai Pra et al., 1996) is
$\mathcal{E}^{\mathbb{P}}\left\{\psi\right\}=\sup_{\mathbb{P}^{\gamma}}\left\{\int\psi d\mathbb{P}^{\gamma}-D_{\text{KL}}\left(\mathbb{P}^{\gamma}\|\mathbb{P}\right)\right\}.$ (3.3)
@@
Dai Pra et al. also provide an analytical formula that the optimal Radon-Nikodym derivative $\frac{d\mathbb{P}^{\gamma^{*}}}{d\mathbb{P}}$ must satisfy:
$\frac{d\mathbb{P}^{\gamma^{*}}}{d\mathbb{P}}=\frac{e^{\Psi}}{\mathbf{E}^{\mathbb{P}}\left[e^{\Psi}\right]},$ (3.4)
under suitable conditions presented in Dai Pra et al. (1996).